OpenStudy (anonymous):

HELP! log(base 25) (6x+25) +log (base 5) (x+25)=5

OpenStudy (anonymous):

\[ \log_{25} (6x+25) +\log_{5} (x+25)=5\]

OpenStudy (mathstudent55):

The sum of logs is the log of the product. log(a) + log(b) = log(ab) Do that to the left side

OpenStudy (mathstudent55):

I just noticed, are the logs on the left side different bases, 5 and 25?

OpenStudy (anonymous):

I changed the base and got \[\log_{5} (6x+25)(x+25)^{2} =\log_{5} 5 ^{10}\]

OpenStudy (mathstudent55):

log(base5)(x + 25) = log(base 5)(x + 25). It's already base 5 log(base 25)(6x + 25) = log(base5)?

OpenStudy (campbell_st):

I think the change of base is incorrect... \[\log_{25}(6x +25) = \frac{\log_{5}(6x + 5)}{\log_{5}(25)} = \frac{1}{2} \log_{5}(6x + 5)\]

OpenStudy (mathstudent55):

log(base25)(6x + 25) = (log(base5)(6x + 25))/log(base5)(25)

OpenStudy (mathstudent55):

log25(6x + 25) = log5( (1/2)(6x + 25) )

OpenStudy (campbell_st):

thats wrong... \[\log_{5}(25) = \log_{5}(5^2) = 2\] so the 1/2 is outside the log... \[\frac{1}{2} \log_{5}(6x + 5) = \log_{5}(6x + 5)^{\frac{1}{2}}\]

OpenStudy (mathstudent55):

@campbell_st Right, my mistake. log25(6x + 25) = (1/2)log5(6x + 25)

OpenStudy (anonymous):

but what will I do with the \[\log_{5} (x+25)=\log_{5} 5^{5} ?? \] after the equation you gave?

OpenStudy (mathstudent55):

That you have the same base on the left side for both logs, you can combine them: